OpenStudy (anonymous):

Can someone help thank you. Write an expression in simplest form for the volume of a cylinder with a radius of 4a2 and a height of 5a. Use 3.14 for pi

OpenStudy (anonymous):

v = (3.14) (4a^2)^2 (5a) = 3.14 * 16a^4 * 5a = 251.2 a^5

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

uw